# Note on Klein-Nishina effect in strong-field QED: the case of nonlinear Compton scattering

### v2

@@ I Introduction
One aspect concerns the Klein-Nishina effect in a strong background field. Another issue is related to the Ritus-Narozhny conjecture, arguing that the account of the background field within the Furry picture is restricted to not too intense fields Ritus:1972ky ; Fedotov:2016afw ; Heinzl:2021mji ; Mironov:2020gbi . In doing so we resort to fairly known formalism as a vehicle to elucidate, numerically, links between established descriptions of linear Compton scattering, nonlinear Compton scattering, and photon emission in a constant crossed field. This subject is also of interest for approximations implemented in simulation codes - both for astrophysics applications and laser physics in particular (cf. DiPiazza:2018bfu ; DiPiazza:2017raw ; Gelfer:2022nqy ; Nielsen:2021nuo ; Blackburn:2021cuq ; Dinu:2019pau ; Ilderton:2018nws ; Heinzl:2020ynb ).
@@
The above (differential) cross section is for the leading-order tree-level term of exclusive $2\to 2$ scattering $e^{-}+\gamma\to{e^{-}}^{\prime}+\gamma^{\prime}$ within perturbative QED, organized in powers of $\alpha$ (for higher orders, cf. Lee:2021iid ). It is known LLIV ; Ritus85 to emerge from nonlinear Compton scattering, $e_{A}^{-}\to{e_{A}^{-}}^{\prime}+\gamma^{\prime}$, as leading-order weak-field term, where $e_{A}^{-}$ is for the electron wave function in an external, classical background field $A$ which may refer to a monochromatic laser field in a plane-wave model as the so-called infinite plane-wave approximation (IPA). Having this relation in mind, one can define a suitably normalized probability (henceforth dubbed ”cross section”) which makes this relation explicit:444 See Bragin:2020akq for the definition of “local cross sections” for processes in an external field. Instead of using wave packages, we employ a circularly polarized, monochromatic plane-wave field as a proxy of a mono-energetic photon beam under special circumstances. The intensity parameter is $a_{0}$.
@@
see Acosta:2021iyu ; Kampfer:2020cbx and Ritus85 for the basics. The classical field strength parameter characterizes the background field $a_{0}$, and the meaning of the variable $u$ is as above. The differential cross section depends separately on two variables: the energy variable $p\cdot k/m^{2}$ and intensity parameter $a_{0}$, which can be combined with the so-called “quantum non-linearity” parameter $\chi=a_{0}\,p\cdot k/m^{2}$ and a second one. The total cross section555 The subsequent numerical evaluations uncover a large range of numerical values of the involved parameters. Therefore, the summation of a large number of harmonics, $N=\mathcal{O}(10^{9})$ in Eq. (9), is often required. (Based on properties of the Bessel functions of large arguments and large indices, harmonics up $N=\mathcal{O}(a_{0}^{3})$ should be accounted for.) In accomplishing that goal, we employ the series expansion method referred to as Richardson extrapolation Sidi:2003bnm , which drastically improves the rate of conversions. $\sigma^{\mathrm{IPA}}$ is compared with $\sigma^{\mathrm{KN}}$ in Fig. 1-left. In addition to the KN effect, the external field causes a further suppression, increasing with $a_{0}$. In line with the above-quoted argument, one can expect that, according to unitarity. In other words, the exclusive decay of an electron state, $e_{A}^{-}\to{e_{A}^{-}}^{\prime}+\gamma^{\prime}$, quantified by the total cross section, becomes more and more reduced with increasing environmental field $A$. The degree of suppression depends on the energy, at a given field strength $A$. This is our executive summary, which is put subsequently into various perspectives in a differential manner. As far as we know, this knowledge appears to be relatively uncommon.
@@
Sometimes, it is claimed in the literature that large values of $\chi$ would refer to high intensities, and the differential probability behind Eq. (9) turns in the expression for one-photon emission in a constant crossed field (CCF). Transferred to “cross section”, linked in an ad hoc manner to Eq. (9), with some seemingly abuse of relations of $\chi$ and $a_{0}$ and energy (see however King:2014wfa for a rationale and discussion in Appendix B), one has  [^6] The locally constant field approximation (LCFA) uses the rate $\frac{dR_{\mathrm{LCFA}}}{d\mathfrak{s}}=\frac{d\sigma^{\mathrm{CCF}}_{a_{0}\to a_{0}g(\phi)}}{d\mathfrak{s}}\frac{a_{0}^{2}g(\phi)^{2}}{4\pi\alpha m^{2}}$, related to CCF Eqs. (14 - 16), where $\mathfrak{s}=u/(1+u)$ is the above mentioned lightfront momentum transfer fraction, see Fedotov:2022ely for details. The quantity $g(\phi)$ is the pulse envelope describing the pulse evolution as a function of the invariant phase $\phi$. Analogously, the locally monochromatic approximation (LMA) is related to IPA via $\frac{dR_{\mathrm{LMA}}}{d\mathfrak{s}}=\frac{d\sigma^{\mathrm{IPA}}_{a_{0}\to a_{0}g(\phi)}}{d\mathfrak{s}}\frac{a_{0}^{2}g(\phi)^{2}}{4\pi\alpha m^{2}}$, where in IPA Eqs. (9 - 13) the replacement $a_{0}\to a_{0}g(\phi)$ also applies and $u=s/(1-s)$ is to be used. (Mimicking pulses by $a_{0}\to a_{0}g(\phi)$ has been traded, according to Heinzl:2020ynb for the first time, in Titov:2019kdk in an ad hoc manner.) Both LCFA and LMA are often used approximation schemes, implemented in many simulation codes for laser-particle/plasma interactions. LMA is shown in Heinzl:2020ynb to approach LCFA in the strong-field limit. The rates $R_{\mathrm{LCFA}}$ and $R_{\mathrm{LMA}}$, defined by Eqs. (84) and (98) in Fedotov:2022ely , require the $\mathfrak{s}$ and $\phi$ integrations, see also Eqs. (1, 2) in Ilderton:2018nws . For more details on LCFA and LMA and improvements, see DiPiazza:2018bfu ; DiPiazza:2017raw ; Ilderton:2018nws .
@@ IV The energy dependence
Turning now to the energy dependence we exhibit in Fig. 3 the differential IPA cross section for $u=0.1$ (left panel), $u=1$ (middle panel) and $u=10$ (right panel). For a given value of $a_{0}$, the IPA cross section matches the CCF expression at small energies, while the high-energy tails coincide with KN. The demarcation is again provided by the first harmonic structures depending on $a_{0}$ and $p\cdot k/m^{2}$ at a given value of $u$ via $u=u_{1}=\frac{2}{1+a_{0}^{2}}p\cdot k/m^{2}$. As mentioned above, besides the total cross section, the differential KN cross section also drops with increasing energy $p\cdot k/m^{2}$. On top of that dropping is the suppression of the differential IPA and CCF cross sections at small energies.
@@ V The $\mathbf{a_{0}}$ dependence
To identify further regions of dominating CCF behavior and discriminate against the KN regime, we display in Fig. 4 the differential cross section at $u=0.1$ (right panel), $u=1$ (middle panel) and $u=10$ (right panel) as contour plots over the $a_{0}$-$p\cdot k/m^{2}$ plane. The differential cross section clearly shows that even for such large values of $\chi$ as $10^{4}$ one encounters the KN regime at large energy $p\cdot k/m^{2}>10^{2}$. The KN regime is that where the contour lines are vertical, i.e. independent of $a_{0}$. In contrast, even for small values of $\chi$, e.g. $10^{0}$, the inclined contour lines at small energies, e.g. $p\cdot k/m^{2}=10^{-1}$, point to the CCF regime. The line $u_{1}=u$, at given $u$, is the demarcation, with a structure related to the lowest harmonics. (Appendix C focuses on the harmonic structures.)
@@
The $a_{0}$-$p\cdot k/m^{2}$ plane can be divided in three regions, YKNexp.supp.CCF at $u=const$, all captured by $d\sigma^{\mathrm{IPA}}/du$ at once. Such a survey might help discuss applicability domains of LCFA and LMA which however require the $u$ integration to arrive at total rates, as pointed out in footnote 6. Our consideration suggest that, on a differential level, LMA turns also into LCFA at high intensities.
VI Many roads lead to large $\chi$
@@
Trading CCF as proxy of $\frac{d\sigma^{IPA}}{du}$, the differential cross section can be fitted by $\frac{d\sigma^{\mathrm{CCF}}}{du}|_{u=const}=\mathcal{A}(u)\,\chi^{-x}$ for $u\in[0.1,20]$ in the asymptotic region, i.e. at large $\chi$. Various trajectories of going to large $\chi$ in the $a_{0}$ vs. $p\cdot k/m^{2}$ plane are exhibited in Fig. 5-left, for definiteness starting at $a_{0}=1$ and $p\cdot k/m^{2}=1$. We find for a path along $a_{0}=const$ (black line): $x=1/3$, along $a_{0}=(p\cdot k/m^{2})^{\beta}$: $x=2/3$ for $\beta=1/2$ (red line), $x=5/6$ for $\beta=1$ (blue line), $x=1$ for $\beta=2$ (green line), and along $p\cdot k/m^{2}=const$ (magenta line): $x=4/3$. That is, $x$ depends strongly on the path of going to large $\chi$. For a path $a_{0}=\mathfrak{z}\left(\frac{p\cdot k}{m^{2}}\right)^{\beta}$, one gets the $\chi$ dependence $\frac{d\sigma^{\mathrm{CCF}}}{du}\propto\mathfrak{z}^{-1/(1+\beta)}\chi^{-1/3-\beta/(1+\beta)}$. The large-$\chi$ limits considered in Ilderton:2019kqp ; Podszus:2018hnz are contained as special cases.
@@
VII Discussion of dependencies and regimes
@@
It comes as some surprise that $d\sigma^{\mathrm{IPA}}_{\mathrm{circ.\,pol.}}\to d\sigma^{\mathrm{CCF}}$, since circular polarization (circ. pol.) means a rotating electric field, i.e. one would expect a closer correspondence of $d\sigma^{\mathrm{IPA}}_{\mathrm{lin.\,pol.}}$ and $d\sigma^{\mathrm{CCF}}$ since CCF means fixed (space-time independent) electric ($E$) and magnetic ($B$) fields, which can be aligned with the linearly polarized (lin. pol.) wave fields. In Appendix B, this issue is considered, as also the meaning of $a_{0}$ w.r.t. CCF.
@@
VIII Summary
Inspired by the ongoing discussion of the Ritus-Narozhny conjecture concerning the validity of the Furry picture for processes at and in external fields, we analyze numerical probabilities of linear and nonlinear Compton scattering. To establish a link to the Klein-Nishina cross section for linear Compton scattering, we chose suitably normalized probabilities as “cross sections”, mainly based on the long-known relation of the weak-field limit of the nonlinear Compton scattering matching the Klein-Nishina expression. Choosing furthermore the monochromatic plane wave model as a proxy of a mono-energetic photon beam, we show with established expressions [^9] Since we envisage a clear cut picture with well-defined precise basics, we resort to the simplest (quantum) formulas describing exclusive one-photon emission. Thus, such a compendium ignores much of the refinements elaborated during the last years, e.g. w.r.t. pulse shapes (implying bandwidth, ponderomotive and interference effects), laser focusing, and the many useful approximation schemes and simulation codes etc., including the efforts of benchmarks. By exhibiting a few relevant plots, we want to display systematically in an intriguing manner the interplay of Klein-Nishina (i.e. linear Compton) and constant-crossed fields to nonlinear Compton and identify - purely numerically - the overlap regions. For a comprehensive survey of modifications of the IPA model, see Fedotov:2022ely . In essence, the parameter space is noticeably enlarged requiring a multitude of specialized considerations of various distinctions. for the one-photon emission in pane waves and constant crossed field how and where the so-defined cross sections coincide. We do our analysis differentially with a Lorentz invariant quantity - the Ritus variable $u$ instead of frequency or/and scattering angle of the produced photon emitted by an electron moving in the external field or scattering at a photon beam. We identify those regions in the three-dimensional parameter space $u$ (for the emitted photon) – $a_{0}$ (external field intensity) – $p\cdot k/m^{2}$ (energy), where the constant crossed field approach is trustable. Since the differential cross section matches inevitably the Klein-Nishina cross section at certain (small) values of $u$, the total cross section contains also such a contribution. We argue that in discussing the Ritus-Narozhny conjecture one should resort to such a differential consideration. It could be that the tails of the Klein-Nishina behavior require adequate treatment in simulation codes, far beyond the improved constant cross field approximations.

### v3

@@ I Introduction
One aspect concerns the Klein-Nishina effect in a strong background field. Another issue is related to the Ritus-Narozhny conjecture, arguing that the account of the background field within the Furry picture is restricted to not too intense fields Ritus:1972ky ; Fedotov:2016afw ; Heinzl:2021mji ; Mironov:2020gbi . In doing so we resort to fairly known formalism as a vehicle to elucidate, numerically, links between established descriptions of linear Compton scattering, nonlinear Compton scattering, and photon emission in a constant crossed field. In this context, we not only review well-known results but also extend the comparative analysis of these findings, which are, in part, widely dispersed across highly specialized publications. This subject is also of interest for approximations implemented in simulation codes - both for astrophysics applications and laser physics in particular (cf. DiPiazza:2018bfu ; DiPiazza:2017raw ; Gelfer:2022nqy ; Nielsen:2021nuo ; Blackburn:2021cuq ; Dinu:2019pau ; Ilderton:2018nws ; Heinzl:2020ynb ).
@@
The above (differential) cross section is for the leading-order tree-level term of exclusive $2\to 2$ scattering $e^{-}+\gamma\to{e^{-}}^{\prime}+\gamma^{\prime}$ within perturbative QED, organized in powers of $\alpha$ (for higher orders, cf. Lee:2021iid ). It is known LLIV ; Ritus85 to emerge from nonlinear Compton scattering, $e_{A}^{-}\to{e_{A}^{-}}^{\prime}+\gamma^{\prime}$, as leading-order weak-field term, where $e_{A}^{-}$ is for the electron wave function in an external, classical background field $A$ which may refer to a monochromatic laser field in a plane-wave model as the so-called infinite plane-wave approximation (IPA). Having this relation in mind, one can define a suitably normalized probability [^4] The relation of Ritus’ differential probability Ritus85 and the differential cross section is given by Ivanov:2004fi ; Acosta:2021iyu $\mathrm{d}W/\mathrm{d}u=a_{0}^{2}m^{2}(p\cdot k)/(e^{2}q^{0})\,\mathrm{d}\sigma/\mathrm{d}u$, where $q^{0}$ is the zeroth component of the effective momentum of the incoming electron. (henceforth dubbed ”cross section”) which makes this relation explicit:555 See Bragin:2020akq for the definition of “local cross sections” for processes in an external field. Instead of using wave packages, we employ a circularly polarized, monochromatic plane-wave field as a proxy of a mono-energetic photon beam under special circumstances. The intensity parameter is $a_{0}$.
@@
see Acosta:2021iyu ; Kampfer:2020cbx and Ritus85 for the basics. The classical field strength parameter characterizes the background field $a_{0}$, and the meaning of the variable $u$ is as above. The differential cross section depends separately on two variables: the energy variable $p\cdot k/m^{2}$ and intensity parameter $a_{0}$, which can be combined with the so-called “quantum non-linearity” parameter $\chi=a_{0}\,p\cdot k/m^{2}$ and a second one. The total cross section666 The subsequent numerical evaluations uncover a large range of numerical values of the involved parameters. Therefore, the summation of a large number of harmonics, $N=\mathcal{O}(10^{9})$ in Eq. (9), is often required. (Based on properties of the Bessel functions of large arguments and large indices, harmonics up $N=\mathcal{O}(a_{0}^{3})$ should be accounted for.) In accomplishing that goal, we employ the series expansion method referred to as Richardson extrapolation Sidi:2003bnm , which drastically improves the rate of convergence. $\sigma^{\mathrm{IPA}}$ is compared with $\sigma^{\mathrm{KN}}$ in Fig. 1-left. In addition to the KN effect, the external field causes a further suppression, increasing with $a_{0}$. In line with the above-quoted argument, one can expect that, according to unitarity. In other words, the exclusive decay of an electron state, $e_{A}^{-}\to{e_{A}^{-}}^{\prime}+\gamma^{\prime}$, quantified by the total cross section, becomes more and more reduced with increasing environmental field $A$. The degree of suppression depends on the energy, at a given field strength $A$. This is our executive summary, which is put subsequently into various perspectives in a differential manner. As far as we know, this knowledge appears to be relatively uncommon.
@@
At large values of $a_{0}$, the differential probability behind Eq. (9) turns in the expression for one-photon emission in a constant crossed field (CCF). Transferred to “cross section”, linked in an ad hoc manner to Eq. (9), with some seemingly abuse of relations of $\chi$ and $a_{0}$ and energy (see however King:2014wfa for a rationale and discussion in Appendix B), one has  [^7] The locally constant field approximation (LCFA) uses the rate $\frac{dR_{\mathrm{LCFA}}}{d\mathfrak{s}}=\frac{d\sigma^{\mathrm{CCF}}_{a_{0}\to a_{0}g(\phi)}}{d\mathfrak{s}}\frac{a_{0}^{2}g(\phi)^{2}}{4\pi\alpha m^{2}}$, related to CCF Eqs. (14 - 16), where $\mathfrak{s}=u/(1+u)$ is the above mentioned lightfront momentum transfer fraction, see Fedotov:2022ely for details. The quantity $g(\phi)$ is the pulse envelope describing the pulse evolution as a function of the invariant phase $\phi$. Analogously, the locally monochromatic approximation (LMA) is related to IPA via $\frac{dR_{\mathrm{LMA}}}{d\mathfrak{s}}=\frac{d\sigma^{\mathrm{IPA}}_{a_{0}\to a_{0}g(\phi)}}{d\mathfrak{s}}\frac{a_{0}^{2}g(\phi)^{2}}{4\pi\alpha m^{2}}$, where in IPA Eqs. (9 - 13) the replacement $a_{0}\to a_{0}g(\phi)$ also applies and $u=s/(1-s)$ is to be used. (Mimicking pulses by $a_{0}\to a_{0}g(\phi)$ has been traded, according to Heinzl:2020ynb for the first time, in Titov:2019kdk in an ad hoc manner.) Both LCFA and LMA are often used approximation schemes, implemented in many simulation codes for laser-particle/plasma interactions. LMA is shown in Heinzl:2020ynb to approach LCFA in the strong-field limit. The rates $R_{\mathrm{LCFA}}$ and $R_{\mathrm{LMA}}$, defined by Eqs. (84) and (98) in Fedotov:2022ely , require the $\mathfrak{s}$ and $\phi$ integrations, see also Eqs. (1, 2) in Ilderton:2018nws . For more details on LCFA and LMA and improvements, see DiPiazza:2018bfu ; DiPiazza:2017raw ; Ilderton:2018nws .
@@ IV The energy dependence
Turning now to the energy dependence we exhibit in Fig. 3 the differential IPA cross section for $u=0.1$ (top panel), $u=1$ (left bottom panel) and $u=10$ (right bottom panel). For a given value of $a_{0}$, the IPA cross section matches the CCF expression at small energies, while the high-energy tails coincide with KN. The demarcation is again provided by the first harmonic structures depending on $a_{0}$ and $p\cdot k/m^{2}$ at a given value of $u$ via $u=u_{1}=\frac{2}{1+a_{0}^{2}}p\cdot k/m^{2}$. As mentioned above, besides the total cross section, the differential KN cross section also drops with increasing energy $p\cdot k/m^{2}$. On top of that dropping is the suppression of the differential IPA and CCF cross sections at small energies.
@@ V The $\mathbf{a_{0}}$ dependence
To identify further regions of dominating CCF behavior and discriminate against the KN regime, we display in Fig. 4 the differential cross section at $u=0.1$ (top panel), $u=1$ (left bottom panel) and $u=10$ (right bottom panel) as contour plots over the $a_{0}$-$p\cdot k/m^{2}$ plane. The differential cross section clearly shows that even for such large values of $\chi$ as $10^{4}$ one encounters the KN regime at large energy $p\cdot k/m^{2}>10^{2}$. The KN regime is that where the contour lines are vertical, i.e. independent of $a_{0}$. In contrast, even for small values of $\chi$, e.g. $10^{0}$, the inclined contour lines at small energies, e.g. $p\cdot k/m^{2}=10^{-1}$, point to the CCF regime. The line $u_{1}=u$, at given $u$, is the demarcation, with a structure related to the lowest harmonics. (Appendix C focuses on the harmonic structures.)
@@
The $a_{0}$-$p\cdot k/m^{2}$ plane can be divided in three regions, YKNexp.supp.CCF at $u=const$, all captured by $d\sigma^{\mathrm{IPA}}/du$ at once. Similarly, this was envisaged in Khokonov:2005 , but is here both quantified and expressed differentially in $u$. Such a survey might help discuss the applicability domains of LCFA and LMA which however require the $u$ integration to arrive at total rates, as pointed out in footnote 7. Our consideration suggests that, on a differential level, LMA turns also into LCFA at high intensities.
VI Accuracy analysis for IPA vs CCF
Figure 5: Example of absolute (left) and relative error (right) as a function of $u$ for $p\cdot k/m^{2}=0.1$ and $a_{0}=10$.
Besides being lengthy examined in previous work DiPiazza:2018bfu ; DiPiazza:2017raw ; Ilderton:2018nws ; Heinzl:2020ynb , the validity of the replacement $\mathrm{IPA}\to\mathrm{CCF}$ for given triple $(a_{0},\chi,u)$ can be controlled for the differential cross-section by the condition
$\frac{u}{N_{\mathrm{ctrl}}}\geq\frac{2\chi}{a_{0}(1+a_{0}^{2})}=\frac{p\cdot k}{m^{2}}\frac{2}{1+a_{0}^{2}},$ (22)
where $N_{\mathrm{ctrl}}$ is the number of harmonics included in IPA (cf. Eq. (9)). Technically, the usage of CCF (instead of IPA), e.g. in a simulation code, for each triple $(a_{0},\chi,u)$ where (22) is fulfilled, introduces an approximation error controlled by $N_{\mathrm{ctrl}}$.
In Fig. 5-left, the absolute approximation error $\mathrm{abs.err.}:=\left|d\sigma^{\mathrm{IPA}}/du-\sigma^{\mathrm{CCF}}/du\right|$ is exemplarily exhibited. It decreases for $u\geq u_{\mathrm{th}}=\frac{p\cdot k}{m^{2}}\frac{2N_{\mathrm{ctrl}}}{1+a_{0}^{2}}$. This approves again the statement that CCF becomes increasingly better for larger $u$. Accordingly, in Fig. 5-right, this is quantified using the relative approximation error $\mathrm{rel.err.}=\frac{\mathrm{abs.err.}}{d\sigma^{\mathrm{IPA}}/du}$. For instance, including only $10$ harmonics, i.e., $N=10$, in the IPA (see l.h.s. of the respective dashed curve) and using CCF on the r.h.s. of this dashed curve introduces a maximum error in the single-digit-percent range and improves [^10] The increasing relative error in Fig. 5-right for $u>1$ has no physical meaning and results from the imprecise numerical cancelations. again for larger $u$.
Consequently, for a given $N_{\mathrm{ctrl}}>5$, the maximum approximation error is given as the value at the respective lower boundary for $u$, i.e. the error value for $(a_{0},\chi,u)$ which solved the equality of (22). Fig. 6-left shows, that, for $a_{0}>1$, the maximum relative approximation error decreases with increasing $N_{\mathrm{ctrl}}$. Furthermore, it is revealed that the error for a given $N_{\mathrm{ctrl}}>5$ does not change drastically, besides numerical cancelations, for a wide range of $a_{0}>1$, which shows the universality of the control parameter $N_{\mathrm{ctrl}}$. This can also be observed in the contour plots in Fig. 4, where a straight line gives the lower boundary of the CCF area in the transition to the IPA. As already discussed in Ilderton:2018nws , in the realm of the local constant-field approximation, the transition from IPA and CCF can be derived analytically by expanding $\mathrm{d}\sigma^{\mathrm{IPA}}/\mathrm{d}u$ in the parameter $\lambda:=\left(\frac{8pk}{m^{2}a_{0}^{2}u}\right)^{\frac{1}{3}}$ and only keep terms linear in $\lambda$. Then, the replacement of IPA with CCF is valid [^11] Similarly, in DiPiazza:2018bfu ; DiPiazza:2017raw , the parameter $\eta_{\mathrm{LCFA}}:=\frac{\chi}{a_{0}^{3}u}$ is proposed to check the validity of LCFA, which is applicable if $\eta_{\mathrm{LCFA}}\ll 1$. In this regard, using condition (22) leads to $1\gg 1/N_{\mathrm{ctrl}}\geq 2\chi/(a_{0}(1+a_{0}^{2})u)\approx\chi/(a_{0}^{3}u)=\eta_{\mathrm{LCFA}}$ for $N_{\mathrm{ctrl}}>1$ and large $a_{0}\gg 1$. if $\lambda\ll 1$.
Using the solution of the equality of the condition (22), the expansion parameter $\lambda$ can be numerically related to the maximum relative error. In Fig. 6-right, we observe that even if $\lambda$ is relatively large, say $\sim 0.7$, the introduced maximum error by using CCF instead of IPA is still in the single-digit percent range. Therefore, by using the condition (22) and especially the control parameter $N_{\mathrm{ctrl}}$, one can quantify, what the $\ll$” in $\lambda\ll 1$ actually means.
Figure 6: Example of the maximum relative error as a function of $N_{\mathrm{ctrl}}$ (left) and $\lambda$ (right) for various values of $a_{0}=1,10,100,1000$, and $p\cdot k/m^{2}=0.1$
VII Many roads lead to large $\chi$
@@
Trading CCF as proxy of $\frac{d\sigma^{\mathrm{IPA}}}{du}$, the differential cross section can be fitted by $\frac{d\sigma^{\mathrm{CCF}}}{du}|_{u=const}=\mathcal{A}(u)\,\chi^{-x}$ for $u\in[0.1,20]$ in the asymptotic region, i.e. at large $\chi$. Various trajectories of going to large $\chi$ in the $a_{0}$ vs. $p\cdot k/m^{2}$ plane are exhibited in Fig. 7-left, for definiteness starting at $a_{0}=1$ and $p\cdot k/m^{2}=1$. We find for a path along $a_{0}=const$ (black line): $x=1/3$, along $a_{0}=(p\cdot k/m^{2})^{\beta}$: $x=2/3$ for $\beta=1/2$ (red line), $x=5/6$ for $\beta=1$ (blue line), $x=1$ for $\beta=2$ (green line), and along $p\cdot k/m^{2}=const$ (magenta line): $x=4/3$. That is, $x$ depends strongly on the path of going to large $\chi$. For a path $a_{0}=\mathfrak{z}\left(\frac{p\cdot k}{m^{2}}\right)^{\beta}$, one gets the $\chi$ dependence $\frac{d\sigma^{\mathrm{CCF}}}{du}\propto\mathfrak{z}^{-1/(1+\beta)}\chi^{-1/3-\beta/(1+\beta)}$. The large-$\chi$ limits considered in Ilderton:2019kqp ; Podszus:2018hnz are contained as special cases.
@@
VIII Discussion of dependencies and regimes
@@
Although, it can be analytically proven Ritus85 , it is somewhat surprising that $d\sigma^{\mathrm{IPA}}_{\mathrm{circ.\,pol.}}\to d\sigma^{\mathrm{CCF}}$, since circular polarization (circ. pol.) means a rotating electric field, i.e. one would expect a closer correspondence of $d\sigma^{\mathrm{IPA}}_{\mathrm{lin.\,pol.}}$ and $d\sigma^{\mathrm{CCF}}$ since CCF means fixed (space-time independent) electric ($E$) and magnetic ($B$) fields, which can be aligned with the linearly polarized (lin. pol.) wave fields. In Appendix B, this issue is examined, along with a discussion of the meaning of $a_{0}$ w.r.t. CCF.
@@
IX Summary
Inspired by the ongoing discussion of the Ritus-Narozhny conjecture concerning the validity of the Furry picture for processes at and in external fields, we analyze numerical probabilities of linear and nonlinear Compton scattering. To establish a link to the Klein-Nishina cross section for linear Compton scattering, we chose suitably normalized probabilities as “cross sections”, mainly based on the long-known relation of the weak-field limit of the nonlinear Compton scattering matching the Klein-Nishina expression. Choosing furthermore the monochromatic plane wave model as a proxy of a mono-energetic photon beam, we show with established expressions [^12] Since we envisage a clear-cut picture with well-defined precise basics, we resort to the simplest (quantum) formulas describing exclusive one-photon emission. Thus, such a compendium ignores much of the refinements elaborated during the last years, e.g. w.r.t. pulse shapes (implying bandwidth, ponderomotive and interference effects), laser focusing, and the many useful approximation schemes and simulation codes, etc., including the efforts of benchmarks. By exhibiting a few relevant plots, we want to display systematically in an intriguing manner the interplay of Klein-Nishina (i.e. linear Compton) and constant-crossed fields to nonlinear Compton and identify - purely numerically - the overlap regions. For a comprehensive survey of modifications of the IPA model, see Fedotov:2022ely . In essence, the parameter space is noticeably enlarged requiring a multitude of specialized considerations of various distinctions. for the one-photon emission in pane waves and constant crossed field how and where the so-defined cross sections coincide. We do our analysis differentially with a Lorentz invariant quantity - the Ritus variable $u$ instead of frequency or/and scattering angle of the produced photon emitted by an electron moving in the external field or scattering at a photon beam. We identify those regions in the three-dimensional parameter space $u$ (for the emitted photon) – $a_{0}$ (external field intensity) – $p\cdot k/m^{2}$ (energy), where the constant crossed field approach is trustable. Since the differential cross section matches inevitably the Klein-Nishina cross section at certain (small) values of $u$, the total cross section contains also such a contribution. We argue that in discussing the Ritus-Narozhny conjecture one should resort to such a differential consideration. It could be that the tails of the Klein-Nishina behavior require adequate treatment in simulation codes, far beyond the improved constant cross field approximations.